A high voltage power transmission line transmitting 100MW of electrical power is operating at a voltage of 288KV. If the line loses 843KW what is its resistance (give your answer in Ohms)?
Answer #1

use,
P = V*I
100*10^6 = 288*10^3*I
I = 347.2 A

power loss = I^2R
843*10^3 = (347.2)^2*R
R = 6.993 ohm
= 7.00 ohm

Answer: 7.00 ohm